OpenStudy (anonymous):

Lines C and D are represented by the equations given below: Line C: y = x + 8 Line D: y = 3x + 2 Which of these shows the solution to the system of equations and explains why? (3, 11), because the point does not lie on any axis (3, 11), because both lines pass through this point (4, 14), because one of the lines passes through this point (4, 14), because the point lies between the two axes

Vocaloid (vocaloid):

set the equations equal to each other x + 8 = 3x + 2 what is x?

OpenStudy (anonymous):

4? o-o

OpenStudy (dan815):

no try again

OpenStudy (anonymous):

3!!! '.'

OpenStudy (anonymous):

wait why is it not 4?

OpenStudy (dan815):

what is 4+8

OpenStudy (anonymous):

8 divided by 2 is 4

OpenStudy (dan815):

what is 3*4 + 2 what is 4+8 4+8 is not equal to 3*4+2 but 3+8=3*3+2 x+8=3*x+2 so when x is 3 its satisfied

Vocaloid (vocaloid):

alternative explanation: x + 8 = 3x + 2 first, we subtract 2 from each side x + 6 = 3x then we subtract x from each side 6 = 2x then we divide by 2 x = 3 does that make sense?